OpenStudy (haleyelizabeth2017):

If cos theta=0.8 and 0

jimthompson5910 (jim_thompson5910):

hint: \[\Large \sin(2\theta) = 2*\sin(\theta)*\cos(\theta)\]

jimthompson5910 (jim_thompson5910):

you'll first need to find the value of sin(theta) before you can use that identity

OpenStudy (abdullahm):

Hint: \(\Large cos ~\theta = \frac{adjacent}{opposite}= 0.8 = \frac{4}{5}\)